OpenStudy (anonymous):

HELLLLLPPPP?(:

OpenStudy (anonymous):

if \[\log_{b}3=0.23 an \log_{b}8=0.71, find \log_{b}24 \]

OpenStudy (anonymous):

round to the nearest thousanths

OpenStudy (anonymous):

@jim_thompson5910

OpenStudy (anonymous):

@timo86m

OpenStudy (anonymous):

yes i know that but its not the answer

jimthompson5910 (jim_thompson5910):

using what HawkCrimson is saying you do this log(24) = log(8*3) log(24) = log(8) + log(3) log(24) = 0.71 + 0.23 log(24) = 0.94 All logs above are base b

OpenStudy (anonymous):

i got log(24) as 1.380

jimthompson5910 (jim_thompson5910):

that's if the base is 10

jimthompson5910 (jim_thompson5910):

but in this case, the base is b and we don't know what b is

OpenStudy (anonymous):

ohhhh, so rounded what would it be

jimthompson5910 (jim_thompson5910):

just 0.94, it's already rounded so to speak

jimthompson5910 (jim_thompson5910):

if you had more than 3 decimal digits, then you would need to round

OpenStudy (anonymous):

ok(: than k you

jimthompson5910 (jim_thompson5910):

yw